OpenStudy (anonymous):

Evaluate: [(a+b)^2 - 4ab(ab)^1/2]^1/2

OpenStudy (anonymous):

\[(a+b)^2 - 4ab \sqrt{ab}\]

OpenStudy (jlastino):

Are there given values for a and b?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

Its 3/2 or 2

OpenStudy (jlastino):

(a+b)^2 - 4a^(3/2) b^(3/2) ?

OpenStudy (experimentx):

\([(a+b)^2 - 4ab(ab)^{1/2}]^{1/2} \) ????

OpenStudy (anonymous):

@Jlastino is correct. I want the square root of that expression

OpenStudy (anonymous):

\[\Large (a+b)^2-4ab\sqrt{ab}\]\[\Large (a+b)^2 -4\sqrt{a^3b^3}\]

OpenStudy (anonymous):

UGLY answer ! lol

OpenStudy (anonymous):

@experimentX It whole to the power 1/2 not -1/2

OpenStudy (anonymous):

@Kreshnik I WANT THE SQUARE ROOT OF THAT EXPRESSION

OpenStudy (anonymous):

isn't square root what I wrote ?

OpenStudy (experimentx):

you know the answer???

OpenStudy (jlastino):

I don't understand the question @_@

OpenStudy (anonymous):

@experimentX I don't know the answer!

OpenStudy (experimentx):

maybe in some ,,, series form??

OpenStudy (jlastino):

The expansion is too long @_@

OpenStudy (anonymous):

The expansion is ugly!

OpenStudy (anonymous):

Anyone?

OpenStudy (anonymous):

experiment x is correct

OpenStudy (anonymous):

What do you mean? Rohangrr

OpenStudy (experimentx):

I can't seem to figure out ... any other way.

OpenStudy (anonymous):

Me too.That's why I posted it here.

OpenStudy (experimentx):

looks like that is the simplest form it could possibly have ... cant think of anything except binomial expansion. BYW was it some part of some other original problem??

OpenStudy (anonymous):

No,my teacher gave it to me this morning and asked if I could figure it out in <30 seconds! The only answer I could give was that for a not equal to b and a>1and b>1 it has no real roots.

OpenStudy (experimentx):

lol if it were ab instead of (ab)^3/2 ... i've could have told in 5 seconds,

OpenStudy (anonymous):

It would have been (a-b).

OpenStudy (experimentx):

if you find out ... please let me know.